AS/AD Model.
AD: P = 48 - 2Y.
AS: P = 8.
PO: Y = 16.
What is the Inflationary Gap
1
Expert's answer
2016-05-15T12:38:02-0400
Find macroeconomic equilibrium, when total demand=total supply
AD=AS
P=48-2Y=8
Y (actual GDP) =20
When actual GDP is above the potential GDP the economy has positive output gap, also called as inflationary gap.
Inflationary Gap = Actual GDP- Potential GDP=20-16=4